# GraphCompNet: A Position-Aware Model for Predicting and Compensating Shape Deviations in 3D Printing

### v1

@@ IV-A Case A: Nesting-Dependent Bar Part Warpage
As shown in Figure 2 (a), the arrangement of the 140 bars within the printing bed and the positions of the sampled data are depicted. The dataset consists of 140 identical bars placed at various positions within the printing bed. Due to the varying thermal profiles across different regions of the bed, shape deviation trends differ significantly. Parts located at the edges of the print chamber, such as Part ID 140 and Part ID 31 in Fig 2 (b), show more significant deviations than those situated at the center of the chamber, such as Part ID 72 and Part ID 24. Figure. 6 provides an example of the observed warpage in “Original CAD versus Scan”, where the color gradient represents the geometric deviation between the original CAD model and the scanned printed part. In this heatmap, blue indicates negative prediction deviation, while red represents positive prediction deviation. It is evident that the center of the bar part experiences higher positive warping, whereas the edges exhibit warping in the opposite direction.
Figure 7: Performance of the proposed DL compensation engine on test data. The heatmap (unit in mm) depicts the geometric distances among various 3D objects, including comparisons between the CAD model and the scanned part, between the CAD model and the compensated CAD, and so forth.
Figure 8: The comparison of C2M signed distances includes the original printed part scan versus the input CAD model, as well as the deformation prediction for the newly compensated part against the input CAD. The latter demonstrates significant improvement.
@@
To train the prediction engine, thirteen bars were randomly selected as training data to capture a range of shape deviation processes within the printer, while the remaining 127 bars were used for evaluating shape deviation prediction and compensation. A point cloud representing the relative position of each part within the print chamber was extracted and input into the proposed architecture. After training the deep learning (DL) prediction engine, its performance was assessed on the test data. An example of the engine’s performance on test data is shown in Figure 6 “Prediction versus Scan”. As shown in the figure, the trained DL prediction engine accurately predicts the deformation of the input CAD model without the need for prior geometric analysis. The predicted part warpage closely matches the scanned printed part, demonstrating that the DL prediction engine effectively minimizes the geometric distance between the predicted and scanned geometries. The DL prediction engine also demonstrates the ability to detect position-dependent shape deviations. For instance, it accurately identifies concave deformation in parts located in the upper region of the print chamber (represented by Part 24 and Part 31 in Fig. 2), and convex deformation in parts situated in the lower region of the printing bed (represented by Part 117 and Part 140).
Next, we evaluate the performance of the DL compensation engine on test data. Figure 7 presents the results of the DL compensation engine on a sample bar part. As shown in Fig. 7 (a) “Original CAD versus Scan”, the geometric distance between the original CAD model and the scanned part reveals significant warping, with negative deviations at the corners and positive deviations at the center. In contrast, the trained compensated part exhibits an opposite, but non-linear trend, as shown in “Original CAD versus Compensated CAD”. The “Original CAD versus Compensated-Predicted CAD” in Fig. 7 (c) demonstrates the application of the DL prediction engine to the compensated CAD, with the goal of aligning the part as closely as possible to the input CAD. The geometric distances between the compensated part and the input CAD are noticeably narrower than those between the original CAD and the scan. This suggests that the DL compensation engine is performing as expected. Additionally, Figure 8 compares the Cloud-to-Mesh (C2M) signed distances between the original printed part scan and the input CAD model, as well as the deformation prediction for the newly compensated part. This comparison quantitatively demonstrates a significant improvement in the geometric accuracy of the compensated part.
The results presented above confirm the theoretical efficacy of the proposed algorithm. However, without physical printing, there is no guarantee that the DL compensation engine has successfully corrected the CAD model in a real-world scenario. Therefore, physical validation through actual printing is essential. To address this, we prepared two sets of printed parts—one set with no compensation and the other with compensation applied. Both sets were printed in the same batch. After printing, we measured the geometric deformation between the uncompensated and compensated part pairs. Over 100 parts from both batches are currently undergoing post-processing and scanning. Figure 9 compares the uncompensated and compensated parts, revealing that while the uncompensated parts exhibit 5.5 mm warpage at the corners, the compensated parts show no noticeable warping.
@@
Figure 11: Comparison of sample part rotated 17 degrees downward along the X-axis, (a) illustrates the difference between the design CAD file and the scanned printed part geometry before applying compensation, (b) shows the difference between the design CAD file and the scanned printed part geometry after applying compensation using our trained prediction and compensation engine.
Compared to our results in Table I, the study by Decker et al. [10] reports an average vertex error reduction of about 44% on a single test part, which aligns with the performance of our model. However, three out of the four parts tested in our study show better compensation results (56% to 66%). Moreover, our approach offers additional benefits by demonstrating greater robustness and accuracy across various print positions and orientations. By accounting for position-dependent variables, which result in different deviation distributions, our method proves to be more versatile and effective in handling a broader range of conditions.
@@ V Conclusions
In this paper, we addressed the critical challenge of achieving high geometric precision in additive manufacturing (AM), focusing on shape deviation modeling and compensation. These deviations, caused by complex interactions among factors such as thermal gradients, and mechanical stresses, hinder AM scalability and efficiency. Existing models often struggle with generalizing compensation strategies for complex 3D geometries and adapting to position-dependent variabilities in batch production. To overcome these limitations, we proposed the GraphCompNet framework, which integrates: 1) point cloud data and DGCNN backbone to capture the irregularities and spatial relationships in complex geometries, enabling the model to learn both local and global patterns of shape deviations. 2) position-specific inputs and a GAN-inspired adversarial training process, to facilitate closed-loop training, enabling real-time compensation and enhancing prediction accuracy under varying print positions.
Our proposed framework reduces the need for extensive experimental validations and computationally expensive models, providing a scalable and precise solution for industrial AM. Through extensive experimental validation, our method demonstrated improvements in compensation accuracy, with geometric deviations of the molded fiber parts reduced by 30% to 66%. The ability to account for position-dependent variations in the print chamber resulted in more efficient compensation strategies. Our approach demonstrated comparable performance to, or outperformed, traditional and state-of-the-art models for single-part deviation prediction and compensation, while also showing robustness across different geometries and print conditions.
Future works include exploring various approaches to encode the continuous position coordinates, such as using sinusoidal (sin/cos) encoding or Fourier features encoding, which could potentially enhance model accuracy and generalization. Further testing on diverse printers, printing conditions, and geometries is essential to deploy the trained model in a real-world industrial setting. For instance, Printer A at Company A may operate under different printing parameters than Printer B at Company B, introducing variables that are outside the distribution of the training data. Consequently, fine-tuning on a minimal set of calibration builds may be necessary to ensure inference accuracy. A thorough investigation into the number of calibration parts required for a new printer, a new set of print parameters, or a new geometry would be highly beneficial to guarantee reliable performance in industrial-scale production. Furthermore, attention should be given to the latency of the serving model, encompassing the point cloud sampling from nominal CAD models, model inference, and the conversion of results back into the proposed new design file. These considerations are crucial to ensure the practical applicability and efficiency of the system in real-world manufacturing environments.
In summary, our study emphasizes the potential of using deep learning techniques to improve geometric precision in AM processes by considering both geometry and the position of each part on the printing bed. By overcoming the limitations of traditional methods, our approach offers significant potential to enhance the accuracy and efficiency of shape deviation compensation across the entire printing space, driving innovation in industrial design and production.
@@ Acknowledgments
This material is based upon work supported by the Air Force Office of Scientific Research under award number FA9550-23-1-0739. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the United States Air Force

### v3

@@ IV-A Case A: Nesting-Dependent Bar Part Warpage
As shown in Fig.2(a), 140 identical bar-shaped parts were arranged across the print bed to analyze the impact of spatial variation on shape deviations. Due to uneven thermal gradients, bars near the edges (e.g., Part IDs 140 and 31) exhibited greater deformation than centrally located parts (e.g., Part IDs 72 and 24), as illustrated in Fig.2(b). Fig.6 provides an example of the observed warpage in “Original CAD versus Scan”, where the color gradient represents the geometric deviation between the original CAD model and the scanned printed part. In this heatmap, blue indicates negative prediction deviation, while red represents positive prediction deviation. It is evident that the center of the bar part experiences higher positive warping, whereas the edges exhibit warping in the opposite direction.
@@
To train the prediction engine, 13 bars were randomly selected, while the remaining 127 were reserved for testing. A point cloud encoding each part’s position in the build chamber was used as input to the deep learning (DL) model. Fig.6(b) demonstrates that the trained prediction engine effectively captures part warpage without prior geometric analysis. The predicted part warpage closely matches the scanned printed part, demonstrating that the DL prediction engine effectively minimizes the geometric distance between the predicted and scanned geometries. The DL prediction engine also demonstrates the ability to detect position-dependent shape deviations. For instance, it accurately identifies concave deformation in parts located in the upper region of the print chamber (represented by Part 24 and Part 31 in Fig. 2), and convex deformation in parts situated in the lower region of the printing bed (represented by Part 117 and Part 140).
@@
We next evaluated the DL compensation engine. As shown in Fig.7, the original printed parts exhibit substantial deviations, especially at the corners. In contrast, the compensated CAD shows inverse adjustments, and when passed through the prediction engine (Fig.7(c)), the predicted deviation is significantly reduced, closely aligning with the original CAD.
The results presented above confirm the theoretical efficacy of the proposed algorithm. However, without physical printing, there is no guarantee that the DL compensation engine has successfully corrected the CAD model in a real-world scenario. Therefore, physical validation through actual printing is essential. To validate physical performance, two sets of parts—compensated and uncompensated—were printed in the same batch. Post-print scanning and comparison reveal that while uncompensated parts showed up to 5.5mm warpage, the compensated parts exhibited minimal or no visible distortion (Fig.8), confirming the real-world efficacy of the proposed method.
@@
Figure 10: Comparison of four sample parts in one print run, the top row illustrates the difference between the design CAD file and the scanned printed part geometry before applying compensation, the bottom row shows the difference between the design CAD file and the scanned printed part geometry after applying compensation using our trained prediction and compensation engine.
IV-C Comparison with State-of-the-art
Compared to our results in Table I, the study by Decker et al. [decker2021geometric] reports an average vertex error reduction of about 44% on a single test part, which is broadly comparable to the performance of our model. However, it is important to note that this comparison is not intended as a direct benchmark, as the two studies involve different printing processes (FDM vs. PBF), which may influence deformation behavior. Instead, this reference is included to qualitatively illustrate the broader challenge of geometric compensation in additive manufacturing. In our study, three out of four parts demonstrate improved compensation performance (56% to 66% error reduction), with our model offering further advantages in robustness and accuracy across varying build positions and orientations. By explicitly accounting for position-dependent variables that lead to diverse deformation patterns, our approach shows greater generalizability and effectiveness under a wider range of conditions. We have revised the manuscript to emphasize that such cross-process comparisons should be interpreted with caution.
@@ V Conclusions
This paper addresses the critical challenge of achieving high geometric precision in additive manufacturing (AM), focusing on modeling and compensating for shape deviations driven by thermal gradients, residual stress, and other process-induced distortions. Existing models often lack generalization across complex geometries and fail to account for position-dependent variability in batch production.
To address these limitations, we proposed GraphCompNet, a graph-based framework that integrates (1) a DGCNN backbone to process point cloud data and capture both local and global geometric patterns, and (2) position-specific inputs within an adversarial training setup to enable closed-loop, position-aware compensation. Experimental validation across various geometries showed consistent improvement, with deviation reductions ranging from 30% to 66%, demonstrating strong generalization and competitive or superior performance to existing methods.
The proposed method offers a scalable and efficient alternative to traditional simulation-heavy workflows by reducing the need for geometry-specific retraining or costly experimentation. Its ability to adapt to spatial variations within the print bed further enhances applicability in industrial settings.
Future work includes: Exploring alternative coordinate encodings (e.g., sinusoidal, Fourier) to improve model accuracy and generalization. Evaluating the number of calibration builds needed for adapting to new printers, process parameters, or geometries. Investigating the latency of the full compensation pipeline for deployment in real-time or near-real-time manufacturing. Investigating adaptive or learnable loss weighting schemes to better balance local and global alignment across diverse shapes and point cloud densities.
Importantly, we emphasize the need for future studies to benchmark compensation strategies under consistent hardware, materials, and print settings. Standardized evaluation protocols will be crucial to rigorously compare methods and advance the state-of-the-art in AM deformation compensation.
In summary, this paper addresses the challenge of compensating for shape deviations in complex geometries and location-dependent variations across batch production in additive manufacturing. By leveraging deep learning techniques, specifically graph-based neural networks, our approach models both global and local geometric interactions while accounting for part-specific positioning on the printing bed. This novel compensation framework overcomes the resolution and scalability limitations of traditional methods, offering an effective solution to enhance geometric precision and production efficiency. Our method demonstrates significant potential for practical implementation in large-scale industrial production, advancing precision-driven innovation in AM processes.
@@ Acknowledgments
This material is based upon work supported by the Air Force Office of Scientific Research under award number FA9550-23-1-0739. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the United States Air Force.
VI Biography Section
[Uncaptioned image] Juheon Lee received his PhD degree from the University of Cambridge, UK, in 2016. Since 2019, he has been a research scientist at HP Inc. His research interests are in geometric deep learning physics-informed neural networks and neural tangent kernel theory. (juheon.lee.626@gmail.com).
[Uncaptioned image] Rachel (Lei) Chen is a Machine Learning Research Engineer at HP Inc. She received her M.S. degree in Electrical and Computer Engineering from Duke University in 2019 and her B.S. degree in Electrical Engineering from Korea Advanced Institute of Science and Technology (KAIST) in 2017. She is now devoted to research in applying cutting-edge deep learning algorithms in Additive Manufacturing quality control, acceleration, and Digital Twins. (lei.rachel.chen@gmail.com)
[Uncaptioned image] Juan Carlos Catana , born in 1987 in Puebla, Mexico, is currently working as a Senior ML Engineer on a 3D software team at HP Inc. His primary research interests include graph algorithms, computational geometry, geometric deep learning, and parallel/distributed algorithms. Outside of work, Juan Carlos enjoys running on the streets, trekking in the mountains, and practicing mountain biking (XC). He is passionate about traveling and discovering new places.
[Uncaptioned image] Hui Wang is an associate professor of industrial engineering at the Florida A&M University-Florida State University College of Engineering and a member of the High-Performance Materials Institute (HPMI). His research has been focused on (i) data modeling and analytics to support quality control for manufacturing processes, including small-sample learning under an interconnected environment and optimal knowledge organization for zero-shot learning, and (ii) optimization of manufacturing system design and supply chain. He received his PhD in industrial engineering from the University of South Florida and an MSE in mechanical engineering from the University of Michigan
[Uncaptioned image] Jun Zeng is a Distinguished Technologist at HP Inc. and a principal investigator and research manager leading software research in 3D Printing and Digital Manufacturing for the HP 3D Printing Software group. His publications include a co-edited book on the Computer-aided Design of microfluidics and biochips, a co-authored book on production management of digital printing factories, 50+ peer-reviewed papers, and 42 granted patents and more pending. His academic training includes advanced degrees in mechanical engineering (PhD) and computer science (M.S.), both from Johns Hopkins University. Jun is an ACM member and an IEEE senior member. (jun.zeng@hp.com).